OpenStudy (anonymous):

Let F(x)= 2x^2+x-3 and G(x)=x-1 find F(x)/G(x) and set its domain

OpenStudy (amistre64):

G not equal zero ... is the only restriction

OpenStudy (ivettef365):

now the question is divide the functions

OpenStudy (amistre64):

or factor f, and remove any g in it :)

OpenStudy (anonymous):

Want me to screen shot the question?

OpenStudy (anonymous):

x divided by x is still x correct?

OpenStudy (amistre64):

no ... x/x = 1 as long as x not= 0

OpenStudy (anonymous):

so it would become 2x^2+1-3?

OpenStudy (amistre64):

if g is a factor in f; then it will have to be of them form: (x-1)(2x+k)

OpenStudy (amistre64):

im not sure what process you are trying to do to come up with that sort of division ....

OpenStudy (anonymous):

This is what i am working with...

OpenStudy (amistre64):

lets narrow it down, when does g=0?

OpenStudy (anonymous):

when x is zero?

OpenStudy (amistre64):

lets try that assumption: g = x-1, let x=0 g = 0-1 = -1. since -1 is not zero .. id say that is not correct

OpenStudy (amistre64):

also, none of your options suggest that x=0 needs to be excluded

OpenStudy (anonymous):

im on like a 7th grade level math... so i need a quick tune up.

OpenStudy (amistre64):

when does x-1 = 0 ? this is pretty basic ....

OpenStudy (amistre64):

there exists some number such that when you remove 1 away from it, you are left with nothing .... what is that number?

OpenStudy (anonymous):

the number is 1

OpenStudy (amistre64):

yes :) so we need to make sure that "1" is not in the domain, so that should narrow it down to 2 options

OpenStudy (anonymous):

LIke how?

OpenStudy (amistre64):

read the options ..

OpenStudy (amistre64):

the domain cannot include the number 1 ...

OpenStudy (anonymous):

it cant be B or C

OpenStudy (amistre64):

B or C is the only options that it can be, so we have to narrow it down between them

OpenStudy (anonymous):

But B and C include 1 in the domain...

OpenStudy (anonymous):

OHhh it says except 1 brain fart.

OpenStudy (amistre64):

no, they include the set of all real numbers EXCEPT 1

OpenStudy (anonymous):

<----- dummy

OpenStudy (amistre64):

now i want you to just take a guess ... which one do you think it has to be correct: B or C?

OpenStudy (anonymous):

B because its left in a fraction.

OpenStudy (amistre64):

no, since the top is NOT the same as f, then that means that g would have to have been canceled out and we would have no bottom to speak about

OpenStudy (anonymous):

Oh. so it has to be C

OpenStudy (amistre64):

let f=gk\[\frac fg=\frac{\cancel gk}{\cancel g}=k\]

OpenStudy (amistre64):

so yes, it has to be C

OpenStudy (anonymous):

@amistre64 see i just figured out how to do that problem now they give me this crap..

OpenStudy (anonymous):

Except 1?

OpenStudy (amistre64):

if x=1, 3-1-10 is not zero

OpenStudy (amistre64):

1+4(30) is not zero, so there are 2 values that need to be restriced .. so none of the options are correct

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

5 times 30 is 150.

OpenStudy (amistre64):

d is wrong simply becasue 1 is not a zero of g a and c are wrong becuase they simply dont exclude any values

OpenStudy (anonymous):

its B!!!

OpenStudy (amistre64):

b is the "most" correct option but its missing one exclusion from the domain

OpenStudy (anonymous):

all real numbers except 2 right?

OpenStudy (amistre64):

except for 2 and -5/3

OpenStudy (amistre64):

so B is the "correct" choice, but its not mathically correct

OpenStudy (amistre64):

the limit of f/g exists at -5/3; but the function has no defined value there

OpenStudy (anonymous):

Yeah i had my teacher look at it. i wont out of this stupid Gradpoint all i run in to is problems.

OpenStudy (amistre64):

spose for instance that your GPS says that you should take a road that crosses at a bridge; the GPS is an equivalent setup for the actual road; but the actual road has the bridge out .... you would not want to take the GPS point at the bridge for the actual point of the bridge missing

OpenStudy (amistre64):

f/g is equal to 1/(x-2) at all points EXCEPT for x=2, or x=-5/3

OpenStudy (anonymous):

Amistre i envy you.

OpenStudy (amistre64):

lol, thnx

OpenStudy (anonymous):

No thank you.

OpenStudy (anonymous):

you really are trying to teach me, i just wish i would have paid attention more to the teacher in math class than the girls. now im screwed for life.

OpenStudy (amistre64):

:) id still pick the girls

OpenStudy (anonymous):

lol :)